# QAEncoder: Towards Aligned Representation Learning in Question Answering System

### v1

@@ 3.3.1. Document Fingerprint Strategy
The distinguishability issue arises because incorporating additional query semantics into document representations suppresses their unique characteristics and renders unrelated documents more similar. To address this issue, we further introduce the document fingerprint strategy, which involves three variations that enhance the uniqueness of document representations from different perspectives.
@@
  • Embedding fingerprint - $\text{{QAE}}_{\text{{emb}}}$. The QAEemb strategy manipulates within the embedding space and reintroduces unique identity of the original document, i.e. the document embedding $\mathcal{E}(d)$. Specifically, $\text{QAE}_{emb}$ considers both the cluster center, $\text{QAE}_{\text{base}}$, and the document embedding, $\mathcal{E}(d)$, balancing their contributions using a hyperparameter $\alpha$. The adjusted embedding is formulated as follows:
    (4) $\text{QAE}_{\text{emb}}(d)=(1-\alpha)\cdot\mathcal{E}(d)+\alpha\cdot\text{QAE}_{\text{base}}(d)\approx(1-\alpha)\cdot\mathcal{E}(d)+\alpha\cdot\frac{1}{n}\sum_{i=1}^{n}\mathcal{E}(q_{i}).$
    Due to the linearity of the inner product space, the cosine similarity, i.e. the inner product between the user query $q$ and $\text{QAE}_{\text{emb}}(d)$, is mathematically equivalent to the weighted sum of the similarities between the user query $q$ and both $\mathcal{E}(d)$ and $\text{QAE}_{\text{base}}(d)$, controlled by hyperparameter $\alpha$.
@@
  • Textual fingerprint - $\text{{QAE}}_{\text{{txt}}}$. The $\text{QAE}_{\text{txt}}$ strategy focuses on the textual space and injects the document identity in a more straightforward manner. Let us define the length of text $c$ in textual space as $|c|$. Before embedding, each document $d$ is enriched by concatenating it with predicted queries such that the ratio of the length of the predicted queries and the length of the original document is about $\beta$. Then, the final embedding is derived as the average representation of these enriched documents.
    (5) $d^{*}=\text{concat}(d,\{q_{j}\}_{j=1}^{k}),\quad s.t.\;|\text{concat}(\{q_{j}\}_{j=1}^{k})|\approx\beta|d|.$
    $\text{QAE}_{\text{txt}}(d)=\frac{1}{n}\sum\limits_{i=1}^{n}\mathcal{E}(d_{i}^{{}^{*}})=\frac{1}{n}\sum\limits_{i=1}^{n}\mathcal{E}(\text{concat}(d,\{q_{j}\}_{j=1}^{k})).$
    For each $d^{*}$, $\{q_{j}\}_{j=1}^{k}$ are randomly and independently sampled from the query generator $\mathcal{Q}(\cdot)$. We point out that the hyperparameter $\beta$ determines the proportion of the original document and the potential queries in the concatenated text. When $\beta$ is low, the concatenated text degrades to the original document; when $\beta$ is high, the document information is overwhelmed by the repeated queries, where the repetition also hinders the query matching.
@@
  • Hybrid fingerprint - $\text{{QAE}}_{\text{{hyb}}}$. The hybrid approach, QAEhyb, seeks to combine the benefits of both the embedding and textual fingerprints for more sophisticated and nuanced document representation. Although QAEemb combines the document embedding $\mathcal{E}(d)$ and the cluster center $\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$ through linear interpolation, inherent differences between these embeddings suggest that a simple linear interpolation should be suboptimal. Therefore, we explore the potential of substituting the document embedding $\mathcal{E}(d)$ in Equation 4 with QAEtxt, which fuses the semantics of both documents and queries.
    (6) $\text{QAE}_{\text{hyb}}(d)=(1-\alpha)\cdot\text{QAE}_{\text{txt}}(d)+\alpha\cdot\text{QAE}_{\text{base}}(d).$
    Note that a more straightforward alternative can be $\text{QAE}_{\text{hyb'}}(d)=(1-\alpha)\cdot\mathcal{E}(d)+\alpha\cdot\text{QAE}_{\text{txt}}(d)$, which replaces $\text{QAE}_{\text{base}}$ in Equation 4 with $\text{QAE}_{\text{txt}}$ that integrates more document information. However, this substitution faces insufficient query semantics and limited improvements.
@@
In our implementation, all the calculated embeddings are normalized for standardized cosine similarity measurement. Our experiments confirm $\text{QAE}_{\text{emb}}$ and $\text{QAE}_{\text{hyb}}$ outperform $\text{QAE}_{\text{base}}$ and $\text{QAE}_{\text{txt}}$.
4. Experiments
@@
To rigorously assess the effectiveness of QAEncoder, we employ six well-known datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019), SQuAD (Rajpurkar et al., 2016), ELI5 (Fan et al., 2019), TriviaQA (Joshi et al., 2017), HotPotQA (Yang et al., 2018) and MSMARCO (Nguyen et al., 2016). These datasets are extensively used for evaluating machine reading comprehension and QA systems, providing a solid foundation for performance evaluation. However, due to classical datasets are frequently utilized for pre-training or fine-tuning embedding models, they fall short in objectively assessing and accurately reflecting the generalized alignment capabilities for queries and documents, particularly for rapidly evolving and updated knowledge base. Recognizing this limitation, we further test on two latest news datasets, the Chinese dataset CRUD-RAG (Lyu et al., 2024) and multilingual dataset FIGNEWS (Zaghouani et al., 2024) covering English, Arabic, French, Hindi and Hebrew. Following the previous work, we evaluate on the development subset; the evaluation metrics include mean reciprocal rank (MRR), mean average precision (MAP), and normalized discounted cumulative gain (NDCG), capturing both the recall and ranking capabilities. We leave the result on MSMARCO, TriviaQA, HotPotQA, FIGNEWS(French), FIGNEWS(Hindi), FIGNEWS(Hebrew) datasets and MAP metric in the appendix for space reason. Refer to the appendix for more implementation details.
4.1. Main Results
@@
We mainly compare QAEncoder against the vanilla encoders. Query-centric methods for sparse retrievers are also presented. Our comparison involves the following approaches:
@@
  • Sparse retrievers - BM25 (Robertson and Zaragoza, 2009), Doc2Query (Cheriton, 2019) and DeepImpact (Mallia et al., 2021b).
@@
  • Dense retrievers - The state-of-the-art embedding models such as BGE models (Xiao et al., 2024) by BAAI, E5 models (Wang et al., 2023b) by Microsoft, GTE models (Zhang et al., 2024b) by Alibaba-NLP, Jina models (Günther et al., 2023) by Jina AI; Other well-known models like Contriever models (Izacard et al., 2021) by Facebook Research, BCEmbedding models (NetEase Youdao, 2023) by NetEase Youdao and the popular Text2Vec models (Xu, 2023). These models feature multilingual understanding and task-specific instruction tuning capabilities. We integrate them with QAEncoder to bridge the document-query gap.
4.1.1. Experiments On Classical Datasets
Table 1. Retrieval performance on classical datasets NQ, SQuAD and ELI5. Higher is better, the best one is bolded. Hyperparameters including QAEncoder variants and weight terms $\alpha,\;\beta$ are optimized simultaneously for six classical datasets. - denotes default or null values.
Model Param NQ SQuAD ELI5
MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8
Sparse
BM25 - 17.9 12.8 46.2 48.2 12.6 11.5
Doc2Query - 18.2 13.6 45.5 48.2 12.6 11.8
DeepImpact - 21.8 15.7 47.8 50.0 12.9 11.9
Dense
bge-large-en-v1.5 - 87.0 68.0 76.2 79.9 57.7 55.8
QAEhyb, $\alpha=0.15,\;\beta=0.5$ 88.2 69.1 78.2 81.8 59.1 57.3
multilingual-e5-large - 86.0 66.3 86.2 88.9 39.0 37.6
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 86.1 66.7 84.9 87.9 46.4 43.3
gte-base-en-v1.5 - 86.2 67.6 68.1 72.4 54.5 51.6
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 85.5 67.4 74.8 78.7 57.0 55.1
jina-embeddings-v2-small-en - 82.4 63.3 69.5 73.6 54.3 51.6
QAEhyb, $\alpha=0.15,\;\beta=0.5$ 83.3 64.1 72.4 76.2 53.8 51.6
contriever - 78.8 60.8 64.8 69.4 51.3 49.8
QAEemb, $\alpha=0.45$ 84.0 65.7 74.9 78.9 55.7 54.4
mcontriever - 52.1 37.6 49.3 54.9 43.0 40.6
QAEhyb, $\alpha=0.45,\;\beta=0.75$ 61.4 45.9 64.7 69.5 51.0 48.3
bce-embedding-base-v1 - 74.4 55.5 77.2 81.3 47.3 45.0
QAEemb, $\alpha=0.3$ 76.4 56.8 77.1 81.1 50.1 48.4
text2vec-base-multilingual - 53.8 36.6 40.9 45.6 38.2 34.7
QAEhyb, $\alpha=0.6,\;\beta=0.5$ 66.6 47.2 56.2 60.6 43.4 40.8
As shown in Table 1, for sparse retrievers, Doc2Query exhibits some improvements over the standard BM25 method. However, the improvements are not as significant as DeepImpact. Conversely, the DeepImpact method, which builds on Doc2Query and incorporates BERT for term weight assignment, achieves better performance across all metrics and datasets. This not only suggests that non-neural retrievers have limited alignment capability, but also highlights the importance of weight adjustment of document and query information.
Compared with sparse retrievers, dense retrievers generally outperform significantly. For the state-of-the-art embedding models such as BGE, E5, and GTE, integrating them with QAEncoder can result in considerable performance enhancements, particularly for rare or unseen datasets. For instance, the multilingual-e5-large model witnesses an MRR increase from 39.0 to 46.4 on the ELI5 dataset, due to the ELI5 dataset is heavily down-sampled during the fine-tuning; while the gte-base-en-v1.5 model improves its MRR from 68.1 to 74.8 on the SQuAD dataset, due to the SQuAD dataset is not included in the fine-tuning data. It suggests that even state-of-the-art embedding models suffer limited generalization, while QAEncoder enjoys robust performance boost. For other popular models as Contriever, BCEmbedding, and Text2Vec, QAEncoder significantly improves them due to their poor generalization caused by relatively smaller training datasets. E.g., the Contriever and mContriever models achieve improvements of 10.5 and 15.5 MRR points on the SQuAD dataset respectively.
4.1.2. Experiments On Latest Datasets
In scenarios such as search engines, financial analysis, and news QA, large volumes of new data constantly emerge and are indexed into retrieval bases for accurate and up-to-date response. Hence, the alignment capability for previously unseen user queries and relevant documents is crucial for embedding models in RAG systems. We experiment on the latest news datasets, the multilingual dataset FIGNEWS and the Chinese dataset CRUD-RAG, to avoid data leakage and mimic the real-world scenarios.
@@ 4.1.2. Experiments On Latest Datasets
Table 2. Retrieval performance on the latest datasets FIGNEWS and CRUD-RAG. Higher is better, the best one is bolded. Hyperparameters including QAEncoder variants and weight terms $\alpha,\;\beta$ are optimized simultaneously for all lastest datasets. - denotes default or null values.
Model Param FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese)
MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8
bge-m3 - 74.3 78.2 77.7 80.5 86.8 88.9
QAEtxt, $\beta=1.5$ 77.1 80.8 80 82.8 89 90.8
multilingual-e5-small - 70.9 74.7 74 77.1 81.7 84.2
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 73.5 77.5 78.8 81.2 89.1 90.6
multilingual-e5-base - 74.7 77.9 72.2 75.7 86.1 88.1
QAEemb, $\alpha=0.3$ 77.4 80.7 77.1 80 88.5 90.5
multilingual-e5-large - 73.8 77.5 76.6 79.9 85.7 88.1
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 77 80.4 82.1 84.9 89.1 91.1
gte-multilingual-base - 65.3 69.9 73.2 76.2 82.9 85.6
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 75.3 79 76.2 79 85.5 88
mcontriever - 32.8 36.3 40.1 44.2 71.7 75.8
QAEhyb, $\alpha=0.45,\;\beta=1.25$ 61.2 65.4 68.2 71.9 88.7 90.5
bce-embedding-base-v1 - 58.9 63.3 - - 76.8 80.3
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 66.6 70.6 - - 85.9 88.1
text2vec-base-multilingual - 38.5 43 27.7 31.4 12.1 13.7
QAEemb, $\alpha=0.75$ 55.2 59.3 51.3 54.8 55.3 58.6

### v3

@@ 3.2.1 Document Fingerprint Strategies
The distinguishability issue arises because incorporating too much query semantics into document representations suppresses their unique characteristics, and renders unrelated documents more similar. To address this issue, we reintroduce the characteristic of documents, intuitively termed document fingerprints, which enhance the uniqueness of QAEbase representations from different perspectives.
@@
QAEbase + Embedding fingerprint = $\text{{QAE}}_{\text{{emb}}}$. The QAEemb strategy manipulates within the embedding space and reintroduces unique identity of the original document, i.e. the document embedding $\mathcal{E}(d)$. Specifically, $\text{QAE}_{\text{emb}}$ considers both the cluster center, $\text{QAE}_{\text{base}}$, and the document embedding, $\mathcal{E}(d)$, balancing their contributions using a hyperparameter $\alpha$. The adjusted embedding is formulated as follows:
$\text{QAE}_{\text{emb}}(d)$ $=(1-\alpha)\cdot\mathcal{E}(d)+\alpha\cdot\text{QAE}_{\text{base}}(d)$ (2)
$\approx(1-\alpha)\cdot\mathcal{E}(d)+\alpha\cdot\frac{1}{n}\sum_{i=1}^{n}\mathcal{E}(q_{i}).$
@@
QAEbase + Textual fingerprint = $\text{{QAE}}_{\text{{txt}}}$. The $\text{QAE}_{\text{txt}}$ strategy focuses on the textual space and injects the document identity in a more straightforward manner. Let us define the length of text $c$ as $|c|$. Before embedding, each document $d$ is enriched by concatenating it with predicted queries to achieve a length ratio of approximately $\beta$ between the queries and the original document. Then, the final embedding is derived as the average representation of these enriched documents.
$d_{i}^{*}=\text{concat}$ $(d,\{q_{j}\}_{j=1}^{k}),\;\text{s.t.}\;|d_{i}^{*}|\approx(1+\beta)|d|.$ (3)
$\text{QAE}_{\text{txt}}(d)=\frac{1}{n}\sum\limits_{i=1}^{n}\mathcal{E}(d_{i}^{{}^{*}}).$
Here $k$ is dynamically determined by the constraint $|d_{i}^{\*}|\approx(1+\beta)|d|$. For each enriched document $d_{i}^{*}$, we first shuffle predicted queries, and iteratively append them after $d$ until reaching the length limit.
@@
Hybrid fingerprint - $\text{{QAE}}_{\text{{hyb}}}$. The hybrid approach, QAEhyb, seeks to combine the benefits of both QAEemb and QAEtxt. Although QAEemb integrates the document embedding $\mathcal{E}(d)$ and the cluster center $\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$ through linear interpolation, inherent differences between these embeddings suggest that a simple linear interpolation should be suboptimal. Therefore, we explore the potential of substituting the document embedding $\mathcal{E}(d)$ in Equation 2 with QAEtxt, which fuses the semantics of both documents and queries.
$\text{QAE}_{\text{hyb}}(d)=(1-\alpha)\cdot\text{QAE}_{\text{txt}}(d)+\alpha\cdot\text{QAE}_{\text{base}}(d).$ (4)
@@
In our implementation, all calculated embeddings are normalized for standardized cosine similarity search. Our experiments confirm $\text{QAE}_{\text{emb}}$ and $\text{QAE}_{\text{hyb}}$ outperform $\text{QAE}_{\text{base}}$ and $\text{QAE}_{\text{txt}}$.
4 Experiments
Model Method AVG ArguAna CQADups. FEVER MSMARCO SciFact Touche20 TRECC.
Sparse
BM25 - 46.9 31.5 29.9 75.3 22.8 66.5 36.7 65.6
DocT5Query - 49.4 34.9 32.5 71.4 33.8 67.5 34.7 71.3
Dense
dpr - 26.4 17.5 15.3 56.2 17.7 31.8 13.1 33.2
QAEemb, $\alpha=0.45$ 37.1 29.6 22.3 70.9 25.4 40.4 22.9 48.0
contriever - 38.1 37.9 28.4 68.2 20.6 64.9 19.3 27.4
QAEemb, $\alpha=0.45$ 45.8 47.1 33.5 73.1 26.1 70.2 25.2 45.2
contriever-msmarco - 49.0 44.6 34.5 75.8 40.7 67.7 20.4 59.6
QAEhyb, $\alpha=0.3,\;\beta=0.75$ 54.9 53.9 38.2 82.0 39.9 73.6 26.3 70.6
bge-large-en-v1.5 - 58.5 63.5 42.2 87.2 42.5 74.6 24.8 74.8
QAEhyb, $\alpha=0.15,\;\beta=0.5$ 61.8 68.8 45.6 91.5 41.2 78.9 28.1 78.2
multilingual-e5-large - 55.0 54.4 39.7 82.8 43.7 70.4 23.1 71.2
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 58.0 61.1 44.3 82.1 43.0 73.9 26.3 75.1
e5-large-v2 - 52.9 46.4 37.9 82.8 43.5 72.2 20.7 66.5
QAEhyb, $\alpha=0.3,\;\beta=1.0$ 57.0 55.1 41.2 86.5 42.8 75.3 23.8 74.2
gte-base-en-v1.5 - 59.4 63.5 39.5 94.8 42.6 76.8 25.2 73.1
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 62.2 68.2 43.7 94.2 41.9 80.3 29.3 77.5
jina-embeddings-v2-small-en - 48.9 46.7 38.0 68.0 37.3 63.9 23.5 65.2
QAEhyb, $\alpha=0.15,\;\beta=0.5$ 54.2 55.3 42.3 74.3 39.5 67.2 27.2 73.6
Table 1: Retrieval performance on seven BEIR benchmarks (NDCG@10). Hyperparameters including QAEncoder variants and weight terms $\alpha,\;\beta$ are optimized simultaneously for all datasets in BEIR. See Table 7 for the full table.
@@ 4 Experiments
To rigorously assess the effectiveness of QAEncoder, we employ well-established BEIR benchmark (Thakur et al., 2021), which contains fifteen publicly available datasets to evaluate the general retrieval performance across diverse domains. However, classical datasets are frequently utilized for pre-training or fine-tuning embedding models  [^1] Please see the fine-tuning data of bge-m3 and bge-multilingual-gemma2-9b., involving ArguAna, FEVER, FiQA, HotpotQA, MSMARCO, NQ, SciDocs and so on (all from BEIR). Hence, classical datasets gradually fall short of objectively reflecting the generalized alignment capabilities for state-of-the-art models, particularly in a rapidly evolving and updated knowledge base.
Recognizing this limitation, we further test on two latest news datasets, the Chinese dataset CRUD-RAG (Lyu et al., 2024) and the multilingual dataset FIGNEWS (Zaghouani et al., 2024) covering English, Arabic, French, Hindi and Hebrew. For BEIR benchmark, we report NDCG@10 as the evaluation metric like previous works (Thakur et al., 2021; Wang et al., 2021; Meng et al., 2022). For the latest datasets, we report both MRR@10 and NDCG@10 metrics, capturing both recall and ranking capabilities. We adopt the cheapest GPT-4o-mini as the main query generator, see Appendix B for query prediction pipeline and cost details.
Hyperparameter Setting.
For the hyperparameters $\alpha$ and $\beta$, we define the following search spaces: $\alpha\in\{0.0,0.15,0.3,0.45,0.6,0.75,0.9\}$ and $\beta\in\{0.25,0.5,0.75,1.0,1.25,1.5\}$. We adopt grid search for QAEhyb. See Appendix H for more hyperparameter selection consideration.
For space reasons, we leave implementation details in Appendix D, the results on complete BEIR benchmark and FIGNEWS datasets in French, Hindi, and Hebrew in Appendix J.
4.1 Main Results
@@
We mainly compare QAEncoder against the vanilla encoders, i.e. the backbones. Query-centric methods for sparse retrievers are also presented. Our comparison involves the following approaches:
@@
Sparse retrievers - BM25 (Robertson and Zaragoza, 2009) and DocT5Query (Cheriton, 2019), an improved version of Doc2Query.
@@
Dense retrievers - The state-of-the-art embedding models such as BGE models (Xiao et al., 2024) by BAAI, E5 models (Wang et al., 2023a) by Microsoft, GTE models (Zhang et al., 2024b) by Alibaba-NLP, Jina models (Günther et al., 2023) by Jina AI; Other well-known models like Contriever models (Izacard et al., 2021) by Facebook Research, BCEmbedding models (NetEase Youdao, 2023) by NetEase Youdao and the popular Text2Vec models (Xu, 2023). For more reference, we also include the seminal dense retriever DPR (Karpukhin et al., 2020), and models fine-tuned on Quora (Thakur et al., 2021), a large dataset of question pairs for question-query retrieval. We integrate them with QAEncoder to bridge the document-query gap.
More details can be found in Appendix F.
4.1.1 Performance on Classical Datasets
As shown in Table 1, for sparse retrievers, DocT5Query effectively augments documents with query prediction, and exhibits significant improvements on the BEIR benchmark over the standard BM25. Dense retrievers gradually outperform sparse retrievers with the scaling-up of model parameter and traning data. For the state-of-the-art embedding models such as BGE, E5, and Jina, integrating them with QAEncoder can lead to robust and generalized alignment, particularly for rare or unseen datasets. For instance, the jina-embeddings-v2-small-en model witnesses a NDCG increase from 65.2 to 73.6 on TRECC dataset; the bge-large-en-v1.5 model’s NDCG on SciFact dataset rises from 74.6 to 78.9; the e5-large-v2 model achieves 8.7 NDCG gains on the ArguAna dataset, as the ArguAna dataset is not included in the fine-tuning data (Wang et al., 2022).
Moreover, QAEncoder significantly improves the performance of other well-known models. For example, the contriever model and its fine-tuned version, contriever-msmarco, achieve 17.8 and 11 NDCG gains on TRECC dataset respectively. More data is available in Tab. 7.
Model Method FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese)
MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10
bge-m3 - 74.4 78.7 77.8 80.9 47.5 48.6
QAEtxt, $\beta=1.5$ 77.2 81 80.2 83.1 51.4 52.5
multilingual-e5-small - 71 75.1 74.1 77.4 44.6 46.0
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 74.6 78.5 78.9 81.6 50.6 51.6
multilingual-e5-base - 74.8 78.1 72.3 76 47.0 48.2
QAEemb, $\alpha=0.3$ 77.6 81.3 77.2 80.3 51.2 52.3
multilingual-e5-large - 73.9 77.8 76.7 80.2 46.9 48.3
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 77.1 80.6 82.2 85.1 51.5 52.7
gte-multilingual-base - 65.5 70.4 73.4 76.8 45.3 46.8
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 75.5 79.5 76.2 79.1 49.4 51.0
mcontriever - 32.9 36.7 40.3 44.7 39.2 41.6
QAEhyb, $\alpha=0.45,\;\beta=1.25$ 61.4 65.9 68.3 72.1 51.3 52.4
bce-embedding-base-v1 - 59.1 63.8 - - 42.0 44.0
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 66.8 71.1 - - 49.7 51.0
text2vec-base-multilingual - 38.7 43.6 27.8 31.9 9.7 10.6
QAEemb, $\alpha=0.75$ 55.4 59.9 51.5 55.4 32.1 34.1
quora-distilbert-multilingual - 39.8 44.2 28.2 32.1 21.8 23.5
QAEemb, $\alpha=0.6$ 50.5 54.7 38.7 42.8 34.4 36.8
@@ 4.1.1 Performance on Classical Datasets
Table 2: Retrieval performance on the latest datasets FIGNEWS and CRUD-RAG. Hyperparameters including QAEncoder variants and weight terms $\alpha,\;\beta$ are optimized simultaneously for all lastest datasets. We leave the full table in Tab. 8, and results on monolingual and bilingual embedding models in Tab. 11 for interested readers.
4.1.2 Performance on Latest Datasets